Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

Key: BOLD: Minutes/notes from last week; BLUE: Notes from this week.

News

  • OSRO/RSRO status below; D configuration delayed at least 3 weeks; final date TBD.
  • PASEO meeting: draft presentations due 28 Jun.
  • Staff Travel:
    • Dickman: 28 Jun - 05 Jul (GB)
    • Lazio: 28 Jun - 06 Jul
    • Butler: 01 Jul - 11 Jul
    • van Moorsel: 26 Jul - 30 Jul
  • Visitors:
    • Sonja Vrcic: 21 June -
    • Brent Carlson: 5 July- 16 July
  • Key Dates:
    • PASEO Meeting: 15-16 July; Claire will be in touch regarding talk titles; review of agenda.

Critical Commissioning review/reports

  • Overall status: Working on resolving antenna/IF-based issues in phase and amplitude; pushing ahead in RSRO (low frequency tested;high frequency underway); some CBE node errors causing issues (new high priority).
  • Highest priority development areas:
    • 1. OSRO blockers
      • Michael's suite of tests used to isolate/classify next level of outstanding issues; high frequency monitoring data taken to accompany tests to elaborate on causes; seeing signs of changes in attenuators despite set-and-remember; also seeing time differences in band changes; faster monitoring enabled to study.*
      • Remaining issues: 1) phase jumps: EA08, EA10, EA11, EA14 Note: removed IF designations as there are indications these can change.; 2) amplitude jumps: EA01 (B), EA06 (B), EA10 (A), EA12 (C); 3) loss of scans (queueing of correlator configurations between CBE/CM);
      • Still need to characterize phase transfer; issues with Cyg A imaging.
    • 2. Wide band widths (2 GHz)
      • Tests (1 hour trials) taken on several more projects. Low frequencies look reasonable (see RSRO section).
    • 3. Switched power implementation
      • Awaiting write-up from Michael (Thursday). Delivered and reviewed. Full implementation not possible prior to Rich's departure; more thought required on software side (particularly for the mapping of information into the SDM format of the SYSPOWER table).
    • 4. 3-bit sampler testing
      • Deployed on 3 antennas; 4th coming soon (EA15, EA22, EA27); Going onto EA28. Initial 'fringes' detected (3 antennas).

System Verification/Hardware Commissioning:

Testing and characterization of the array based on the design goals.

Targets

Area Target Priority JIRA Deliver Status
* Switched power gain correction 0 - pending SDM implementation for further work  
* RFI Sweep for all bands 2 - 2010Q3 Pending wide band widths

Reports

  • Antenna: Perley
    • EA06 status (pointing, working rxs); IF B problematic; replaced DTS module; needs work beyond maintenance day. Module replaced; problem fixed.
    • EA20 amplitude modulations
  • Receiver: Perley
    • Ku rxs coming (2) next week.
  • Array: Perley
    • Succinct list of issues
      • Phase jumps: EA04, EA08, EA10, EA11, EA14
      • Amp jumps: EA01 (B), EA06 (B), EA10 (A), EA12 (C); EA20 unreliable
  • Correlator: Rupen
    • 5% zeros/dropped data due to loss of synchronization; working to improve this (Carlson); 1% switching bands; few% switching configurations.
    • Sonja has list of priorities; pushing on CM clearing of configs; queuing and reconfigs, etc

Observing Band Checkout

Encompasses principally core operation at those frequencies followed by detailed calibration strategies, flagging.

Targets

Reports

  • Low Frequency (L/S/C): Momjian
    • News/Data Updates: 10B-187 data review: Baselines to EA02 (E02) show a comb in Stokes LL throughout the observations (for instance, see the attached EA02_SPEC.png). The sky frequency is 1420.40575 MHz, and the bandwidth is 1 MHz. The spectral spikes are separated by 0.125 MHz. (ECDATA-261). Being reviewed by Kerry.
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-38

  • Mid Frequency (X/Ku): Dhawan
    • Status: Operating; users should include one of the canonical flux calibrators in SB.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links

  • High Frequency (K/Ka/Q): Sjouwerman/Claussen
    • Status:
      • Operating; users should include one of the canonical flux calibrators in SB
      • Reference pointing in C or X
      • No observing at Ka < 32 GHz; No longer blocked; SSS/OPT code is consistent with limitations in hardware.
    • Upcoming:
      • Need switched power
        • Develop set of observations to check Tcal/Tsys/SEFD across bands
    • Links: ECSV-30

Observing Mode Support:

Testing and development within broad observing modes distinguished by their science goals (continuum, spectral line, polarization).

Targets

Area Target Priority JIRA Deliver Status
Spectral Line Bandpass stability 1 ECSV-35 2010Q3  
Spectral Line Doppler tracking 1 ECSV-32 2010Q3  
Spectral Line Spectral line setups; calibration procedures recommendations (OSRO/RSRO) 1 ECSV-31 2010Q3  
Polarization On-axis leakage calibration 1 ECSV-25 2010Q3  
Polarization R-L phase calibration 1 ECSV-25 2010Q3  
Polarization Observer Recommendations (OSRO/RSRO) 2 ECSV-25 2010Q3  
Polarization Beam squint calibration 2 ECSV-25 2010Q3  
Polarization Primary beam (off-axis) calibration 2 ECSV-25 2010Q3  
Polarization Polarization calibration service observation plan 3 ECSV-25 2010Q4  
Demo Science SBs for approved EPO projects 1 ECSV-57 2010Q3 Started for a subset

Reports

  • Spectral Line Modes (Ott/Goss): TSPE
    • Testing:
      • Spectral Line general commissioning tests (diff spw widths, diff frequencies, bandpasses, etc): ECSV-92 * Bandpass stability: ECSV-35: Report
    • Links: ECSV-19
  • Polarization Modes (Myers): TPOL
    • Status:
    • Testing
      • Polarization calibration L through Q:
      • C (ECDATA-12): Ready
      • L (ECDATA-42): Ready
      • Combined Low-High Pol checkout test (TPOL0003_sb1378180): Report
      • Post-processing (AIPS): AH1004 data checkout
    • Links: ECSV-29

  • Demo Science (Chandler/Condon): TDEM
    • Status: Begun; list requires prioritization; limited remaining time in D-config
    • Testing
      • Cyg A (Perley): SBs ready; Issues as noted above; will need repeat.
      • Crab Nebula (Frail): SBs submitted; 1 hour trial report (Frail)
    • Links: https://staff.nrao.edu/wiki/bin/view/EVLA/DemoScience

  • Advanced Planetary: TPLA
    • Testing
      • Cassini observations obtained.; Status.
    • Status:
      • Executor: Ready for testing (Butler).
      • SDM: Ephemerides/motion not supported.
  • Advanced: Solar: TSOL
    • Status: Planning/prototype testing ongoing; working with L, Ku (special Rx setups in these two bands). Perley: On sun fringe check with WIDAR.
  • Advanced: Pulsar (Brisken): TPUL
  • Advanced: VLBI (Walker/Brisken): TVLB
  • Advanced: Astrometry (Mioduszewski)

Observing Support:

Testing and development of the end-to-end capability of the array to support observations to the community (Documentation, OPT, Scheduling, Instrument models, Calibrator survey/models DB, Rapid response observing, demo science, post-processing).

Targets

Area Target Priority JIRA Deliver Status
Software CASA development cycle priorities 0 -    
Software CBE averaging/decimation (time/freq) 1 - 2010Q3; vci averaging - honors other-than-1s dumps.  
Software/Calibration Recommended data path for weights in calibration 1 - 2010Q3  
Flux Calibration SBs for Flux calibrator models 1 ECSV-43 2010Q3 Needed for D config (2 freqs/band)

Reports

  • Meta summary
    • Multi-config SBs in regular use; note 'overhead' in configuration changes; Must add 30 seconds to each configuration change - Gustaaf's group will check projects for this until relaxed, in particular, reference pointing must expand to 3 minutes..

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
  wide band modes available only in webtest OST (reconciling status database)

OPT selected limits not possible
Running 20 Hz
18-May CM version

21-May CBE version

    - needs check on short pol'n runs Need decision on large (i.e., RSRO) data set distribution media   Operational for OSRO1/OSRO2
30s overhead with OSRO configuration changes
RSRO beginning

  • Software Systems (Overlord: Butler):
    • Versioning Page: https://staff.nrao.edu/wiki/bin/view/EVLA/EVLASoftwareVersioning Updated. Tracking builds/deployment, release notes, current.
    • CBE store - new lustre file system next Wednesday.
    • OPT (Lorant): new production version today; see release notes for fixes; would like projects from webtest to be submittable; may not be needed as SB # can be accessed. Deployed and in use.
    • Scheduling (Joan):
      • *Truitt reviewing issue list from tests to improve heuristics for project selection/execution.Updated; ready to test.
    • Executor (Barry): Changes this week See release notes for changes (changes for 3-bit, phased array).
    • Corr (CBE/CM; Michael): Sonja in town; focusing on CM queueing issues/configuration clearing initially.
    • Archive (John):
    • Post-processing
      • CASA (Nick):
        • Development plan out the door; being finalized then will post; most priority 1 development items covered; some speed improvements in flagging (being tested by Gustaaf/Joe); proposed change to flagdata interface; needs more thought (asked Miriam to lead a draft set of use cases to refine the interface options); problem if 'source' init casapy script (when adding own tasks) and running casapy - Wes is looking into it; discussion of timing of updates to CASA dependent libraries - can request updates to be judged by Nick.
      • AIPS (Eric): REWAY statistics
    • ASDM
      • Priority 1: SDM total power/switched power table write-up (Rupen). Done.
      • Priority 2: Flag table write-up (Myers/McMullin) Due. Needed for 5 Jul ALMA FBT
      • Priority 3: Ephemeris table resolution (Butler)

  • Flux Calibrator Models (Amy/Rick): TCAL * Links: ECSV-43

  • RSRO: TRSR
    • Status: RSRO Readiness (Cycle1) Master List
      • TRSR0005 (AB1346): 25Jun trial bad (Brogan)
      • TRSR0001 (AB1345): Sanjay, Rick, Michael
      • TRSR0018: Laura
    • Resident Schedule: Posted on home page of twiki now!
      • 01 Jan - 16 May: Adam Deller (AC982: RSRO-4): 300A
      • 31 Jan - 15 Jul: Laura Chomiuk (AS1015/AK726: RSRO-11/RSRO-6): 200C
      • 04 Feb - 01 Jun: Jacqueline van Gorkom (AV318: RSRO-13): 200D
      • 25 Feb - 01 Jun: James Miller-Jones (AM1014: RSRO-9): 200E
      • 02 Mar - 07 Jun: Manuel Aravena (AA330: RSRO-1): 300G
      • 16 Mar - 08 Apr: Dave Green (AB1345: RSRO-2): 300G
      • 20 Apr - 30 Jun: Crystal Brogan (AB1346: RSRO-3): 200B
      • 20 Apr - 30 Jun: Todd Hunter (AB1346: RSRO-3): 200B
      • 30 Apr - 15 Jul: Joseph Lazio (AC982: RSRO-4): 266
      • 07 Jun - 16 Sep: Laura Perez (AC982: RSRO-4): 200E
      • 28 Jun - 05 Oct: Volker Heesen (AH1006: RSRO-5): 200G
      • 05 Jul - 26 Oct: Russ Taylor (AT374: RSRO-12): 300G
      • 21 Jul - 23 Sep: Dave Green (AB1345: RSRO-2): 300G
      • 15 Sep - 15 Dec: Adam Leroy (AL746: RSRO-7): 200C
      • 01 Oct - 01 Jan: Hendrik Linz (AC982: RSRO-4): 200C
      • 01 Jan - 01 Apr: Jane Greaves (AC982: RSRO-4): 300G
      • 02 Jan - 02 Apr: Stuart Corder (AC982: RSRO-4): 300G

    • Current Residents
      • Crystal Brogan (POC: Claire Chandler); area(s): High frequency observing; demo science
        • This week: Review of Ka selfcal issues (with Myers)
      • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line commissioning; OSRO data validation
        • This week: Not available
      • Laura Perez (POC: Claire Chandler); area(s): High frequency observing
        • This week: AC982 prep; RSRO/OSRO pointing analysis (as available);
      • Todd Hunter (POC: Claire Chandler; area(S): science support software interfaces for users and high frequency observing; band/polarization checkout.
      • Joe Lazio (POC: Claire Chandler): Low frequency observing.
        • This week: Phased array tests; rapid response pipeline; RSRO/OSRO pointing tests
    • Past (and future residents)
      • Manuel Aravena (POC: Chris Carilli); area(s): high frequency observing
      • Adam Deller (POC: Michael Rupen); area(s): phased array system design and implementation
      • Dave Green (POC: Sanjay Bhatnagar): area(s): algorithm tests for wideband imaging (currently away).
      • James Miller-Jones (POC: Claire Chandler; area(s): demo science use cases
      • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; spectral line commissioning)

  • Science Support Team
    • McMullin: Scheduling/On-call; test data reviews; review X/C pointing results; PASEO Prep; FLAG table (with Steve)
    • Myers: Review of TPOL0002 with Crystal looking at cal transfer
    • Ott: CASA development planning; work for CDR 8; submit M51 SBs (demo science); helpdesk tickets.
    • Greisen: catch-up on e-mail; AIPS letter; change scaling on cubes (Jy/beam!).
    • Dhawan: data reviews from targeted tests, pointing (plots, reports).
    • Kogan: FRMAP task tutorial
    • Rupen: Switched power writeup; testing to isolate outstanding OSRO blockers; RSRO testing; work with Sonja.
    • Krauss: Reviewing pipelines; developing EVLA doc pages, improve organization, particularly for non-radio astronomers.

Advanced Science Support:

These are the testing/development areas needed to fulfill the full capabilities of the EVLA along with additional observing modes for other branches of the community.

  • Astrometry (Mioduszewski)
    • Status: Wait for A array(?)

AOB

Misc Actions

  • [Rupen]: Develop daily/post-maintenance day checklists for the system. To be completed as part of stress tests.
    • Evolving to post-software/firmware/hardware update suite of tests
    • Stress tests still to come.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Wrobel/Butler]: Make use of API criteria consistent throughout tools (OPT->OST).
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Med Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; understood as a synchronization issue; currently at 5% of the data for 'most' observations. Carlson/Fort at work on mitigating/eliminating.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 28 Jun 2010
Topic revision: r4 - 2010-06-29, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback